### v1

@@ 6 Convergence of the divergence-conforming discretisation
$|||(\boldsymbol{u}-\boldsymbol{u}_{h},p-p_{h},\theta-\theta_{h},\lambda-\lambda_{h})|||^{2}\lesssim|||(\boldsymbol{u}-\tilde{\boldsymbol{u}},p-\tilde{p},\theta-\tilde{\theta},\lambda-\tilde{\lambda})|||^{2}+\sum_{K\in{\mathcal{T}}_{h}}h_{K}^{2}|\boldsymbol{u}-\tilde{\boldsymbol{u}}|_{2,K}^{2}.$
Moreover, if $(\boldsymbol{u},p,\theta,\lambda)\in\boldsymbol{\mathrm{H}}^{k+2}(\Omega)\cap\boldsymbol{\mathrm{H}}\times H^{k+1}(\Omega)\cap Q\times H^{k+2}(\Omega)\cap W\times H^{k+1/2}(\Gamma_{m})\cap Z$, then
$|||(\boldsymbol{u}-\boldsymbol{u}_{h},p-p_{h},\theta-\theta_{h},\lambda-\lambda_{h})|||^{2}\lesssim h^{k+1}.$
@@
$|||(\boldsymbol{u}-\boldsymbol{u}_{h},p-p_{h},\theta-\theta_{h},\lambda-\lambda_{h})|||$ $\leq|||(\boldsymbol{u}-\tilde{\boldsymbol{u}},p-\tilde{p},\theta-\tilde{\theta},\lambda-\tilde{\lambda})|||$
$\qquad+|||(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})|||.$ (43)
Next we derive the bound of $|||(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})|||$. Using Lemma 5.7, for each $(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})\in{\boldsymbol{\mathrm{H}}_{h,0}}\times Q_{h}\times Z_{h}\times W_{h}$, we find $(\boldsymbol{v},q,\tau,\xi)\in{\boldsymbol{\mathrm{H}}_{h,0}}\times Q_{h}\times Z_{h}\times W_{h}$ with
$|||(\boldsymbol{v},q,\tau,\xi)|||\lesssim|||(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})|||,$
@@
$|||(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})|||^{2}\lesssim B(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h};\boldsymbol{v},q,\tau,\xi),$
where
$B(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h};\boldsymbol{v},q,\tau,\xi)=a_{h}(\boldsymbol{u}_{h},\boldsymbol{v})+b_{h}(\boldsymbol{v},(p_{h},\lambda_{h}))+b_{h}(\boldsymbol{u}_{h},(q,\xi))+c(\theta_{h},\tau).$
@@
$|||(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})|||^{2}$ $\lesssim B(\tilde{\boldsymbol{u}},\tilde{p},\tilde{\theta},\tilde{\lambda};\boldsymbol{v},q,\tau,\xi)-B(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h};\boldsymbol{v},q,\tau,\xi)$
@@
$\lesssim B(\tilde{\boldsymbol{u}}-\boldsymbol{u},\tilde{p}-p,\tilde{\theta}-\theta,\tilde{\lambda}-\lambda;\boldsymbol{v},q,\tau,\xi)+\mathcal{R}_{em},$ (44)
where
@@
$\mathcal{R}_{em}:=\widetilde{a}_{h}(\boldsymbol{u},\boldsymbol{u},\boldsymbol{v})-\widetilde{a}_{h}(\boldsymbol{u}_{h},\boldsymbol{u}_{h},\boldsymbol{v})+\widetilde{c}_{h}(\boldsymbol{u},\theta,\tau)-\widetilde{c}_{h}(\boldsymbol{u}_{h},\theta_{h},\tau)+\langle\xi,g(\theta)-g(\theta_{h})\rangle_{\Gamma_{m}}.$
Using the continuity bounds implies that
$|\widetilde{a}_{h}(\boldsymbol{u},\boldsymbol{u},\boldsymbol{v})-\widetilde{a}_{h}(\boldsymbol{u}_{h},\boldsymbol{u}_{h},\boldsymbol{v})|$ $\leq M\big{(}\|(\boldsymbol{u}-\tilde{\boldsymbol{u}})\|_{1,h}+\|(\boldsymbol{u}_{h}-\tilde{\boldsymbol{u}})\|_{1,h}\big{)}\|\boldsymbol{v}\|_{1,h},$ (45a)
$|\widetilde{c}(\boldsymbol{u},\theta,\tau)-\widetilde{c}(\boldsymbol{u},\theta,\tau)|$ $\leq M\big{(}||\theta-\tilde{\theta}||_{1,\Omega}+||\theta_{h}-\tilde{\theta}||_{1,\Omega}+\|(\boldsymbol{u}-\tilde{\boldsymbol{u}})\|_{1,h}+\|(\boldsymbol{u}_{h}-\tilde{\boldsymbol{u}})\|_{1,h}\big{)}||\tau||_{1,\Omega},$ (45b)
$|\langle\xi,g(\theta)-g(\theta_{h})\rangle_{\Gamma_{m}}|$ $\leq L||\xi||_{-1/2,\Gamma_{m,h}}(||(\theta-\tilde{\theta})||_{1,\Omega}+||(\theta_{h}-\tilde{\theta})||_{1,\Omega}),$ (45c)
where $M$ and $L$ are sufficiently small. Combining (6.2) and (6.2) yields that
@@
$|||(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})|||^{2}$ $\lesssim|||(\boldsymbol{u}-\tilde{\boldsymbol{u}},p-\tilde{p},\theta-\tilde{\theta},\lambda-\tilde{\lambda})|||^{2}+\sum_{K\in{\mathcal{T}}_{h}}h_{K}^{2}|\boldsymbol{u}-\tilde{\boldsymbol{u}}|_{2,K}^{2}.$
@@
Substituting the above bound in (6.2) leads to the stated estimate. Using the approximation results given in [9, 37, 38, 25, 27] leads to the second stated result.
@@ 7 Numerical experiments
[…]
@@
To compute $\|\lambda-\lambda_{h}\|_{s,}$ (with $s\in\{-\frac{1}{2},\frac{1}{2}\}$ because we will use Lagrange multipliers in these two spaces) we use the characterisation of $\mathrm{H}^{s}(\Gamma_{\mathrm{m}})$ in terms of the spectral decomposition of the Laplacian operator (see, e.g., [29]). For this, let $R:\mathrm{H}^{1}(\Gamma_{\mathrm{m}})\rightarrow\mathrm{H}^{1}(\Gamma_{\mathrm{m}})$ be the bounded linear operator defined by
@@
and so $\mathrm{H}^{s}(\Gamma_{\mathrm{m}})$ is the closure of the span of $\{r_{i}\}_{i=1}^{\infty}$ in this norm. During the experiments, different values for the stabilisation parameters will be considered in order to capture the convergence of the method.
@@ 7.1 Divergence-conforming test
Table 7.1: Example 7.1. Error history and Newton iteration count for a finite element family of $\mathbb{BDM}_{k+1}-\mathbb{P}_{k}-\mathbb{P}_{k}-\mathbb{P}_{k+1}$, with $k=0,1,$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively, on the unit square domain $\Omega=(0,1)^{2}$. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})-$norm, and $\alpha_{0}=10(k+2)$.
@@
$\boldsymbol{u}(x,y,z)=\left(\begin{aligned} &\sin(\pi z)\,\cos(\pi y)\\ &-\cos(\pi x)\,\sin(\pi z)\\ &\sin(\pi x)\,\cos(\pi y)\end{aligned}\right),\qquad p(x,y,z)=\sin(x^{2}+y^{2}+z^{2}),\qquad\theta(x,y)=e^{-xyz}.$
[…]
@@
222 1.000 2.32e+00 $\star$ 1.73e+00 $\star$ 2.81e+00 $\star$ 1.80e-01 $\star$ 4
4478 0.554 1.11e+00 1.26 1.04e+00 0.87 9.22e-01 1.89 8.82e-02 1.21 4
@@
Table 7.2: Example 7.1. Error history and Newton iteration count for a finite element family of $\mathbb{BDM}_{k+1}-\mathbb{P}_{k}-\mathbb{P}_{k}-\mathbb{P}_{k+1}$, with $k=0,1,$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively, on the unit cube domain $\Omega=(0,1)^{3}$. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})-$norm, and $\alpha_{0}=10(k+2)$.
7.2 Filtration with osmotic effects
$\kappa=4955.144\,J/\mathrm{mol},\quad A_{0}=1.189\cdot 10^{-11}\mathrm{m}\,\mathrm{Pa}^{-1}\mathrm{s}^{-1},\quad\mu_{0}=8.9\cdot 10^{-4}\mathrm{Kg}\,\mathrm{m}^{-1}\mathrm{s}^{-1},$
$\rho_{0}=1027.2\,\mathrm{Kg\,m}^{-3},\quad D_{0}=1.5\cdot 10^{-9}\mathrm{m}^{2}\,\mathrm{s}^{-1},\quad\Delta P=4053000\,\mathrm{Pa}.$
With respect to the boundary conditions on the inlet, we will consider the following:
$\boldsymbol{u}\cdot\boldsymbol{n}=\left(u_{0}-v_{w}\frac{2x}{d}\right)\left(\frac{3}{2}(1-\lambda^{2})\right)\left[1-\frac{\mathrm{Re}}{420}\left(2-7\lambda^{2}-7\lambda^{4}\right)\right],$
$\boldsymbol{u}_{2,h}$
$\theta_{h}$efer to caption
To capture the velocity behaviour at the inlet, as well as the maximum permeate velocity, a free tangential stress $(\boldsymbol{\sigma}\boldsymbol{n})\cdot\boldsymbol{t}=0$ is imposed at $\Gamma_{\mathrm{in}}$. Similar results for the dual membrane channel are obtained if we consider $\boldsymbol{u}_{2}$ as the corresponding velocity component on a Berman flow. The results for the first and second scenarios are depicted in Figure 7.1. For the first case we can see that the velocity near the membrane is affected by the porosity and the transmembrane pressure. In addition, we notice, at the inlet (due to the minimal amount of salt compared to the rest of the membrane), a high flux in the normal direction of the velocity with respect to the membrane. In Figure 7.2 we compare the performance of both channels. It can be seen that the concentration profile and permeate velocity are highly dominated by the transmembrane pressure, irrespective of the choice of inlet profiles. On the other hand, we observe that the concentration profile at the membrane increases as we approach the end of the channel, consequently decreasing the permeate velocity. This is accompanied by a linear pressure drop, which behaves similarly for both scenarios.
7.3 A channel with a spacer
In this test we study the behaviour of the proposed method when an obstacle, serving a a spacer, is considered. More precisely, we consider a channel with the same dimensions as the previous experiment, but with the addition of a spacer in the middle of the channel in a cavity-type configuration. The spacer corresponds to the cross section of a cylinder, i.e., a circle, with diameter $3.6\cdot 10^{-4}$m and tangent to the membrane. The boundary conditions for the spacer are the same as $\Gamma_{\mathrm{wall}}$. To study the behaviour of the method, the velocities to be tested in this experiment are $u_{0}=5\cdot 10^{-2}$m///s and $u_{0}=1.29\cdot 10^{-1}$m///s, and the effect of the salt concentration boundary layer along the membrane is studied.
$|\boldsymbol{u}_{h}|$
$\theta_{h}$efer to caption
Figure 7.3: Velocity streamlines (top panels) and concentration profiles (bottom panels) around the spacer in a cavity-type configuration and inlet velocities $u_{0}=5.0\cdot 10^{-2}$m///s (left) and $u_{0}=1.29\cdot 10^{-1}$m///s. The bottom numbers indicate distance from inlet (in ×$\times 10^{-3}$m).
Appendix A Lagrange multiplier stabilisation for a conforming approximation

### v2

@@ 6 Convergence of the divergence-conforming discretisation
$\lvert\!\lvert\!\lvert(\boldsymbol{u}-\boldsymbol{u}_{h},p-p_{h},\theta-\theta_{h},\lambda-\lambda_{h})\rvert\!\rvert\!\rvert^{2}\lesssim\lvert\!\lvert\!\lvert(\boldsymbol{u}-\tilde{\boldsymbol{u}},p-\tilde{p},\theta-\tilde{\theta},\lambda-\tilde{\lambda})\rvert\!\rvert\!\rvert^{2}+\sum_{K\in{\mathcal{T}}_{h}}h_{K}^{2}|\boldsymbol{u}-\tilde{\boldsymbol{u}}|_{2,K}^{2}.$
Moreover, if $(\boldsymbol{u},p,\theta,\lambda)\in\boldsymbol{\mathrm{H}}^{k+2}(\Omega)\cap\boldsymbol{\mathrm{H}}\times\mathrm{H}^{k+1}(\Omega)\cap Q\times\mathrm{H}^{k+2}(\Omega)\cap W\times\mathrm{H}^{k+1/2}(\Gamma_{m})\cap Z$, then
$\lvert\!\lvert\!\lvert(\boldsymbol{u}-\boldsymbol{u}_{h},p-p_{h},\theta-\theta_{h},\lambda-\lambda_{h})\rvert\!\rvert\!\rvert^{2}\lesssim h^{k+1}.$
@@
[…]
(44) $\qquad+\lvert\!\lvert\!\lvert(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})\rvert\!\rvert\!\rvert$.
Next we derive the bound of $\lvert\!\lvert\!\lvert(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})\rvert\!\rvert\!\rvert$. Using Lemma 5.7, for each $(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})\in{\boldsymbol{\mathrm{H}}_{h,0}}\times Q_{h}\times Z_{h}\times W_{h}$, we find $(\boldsymbol{v},q,\tau,\xi)\in{\boldsymbol{\mathrm{H}}_{h,0}}\times Q_{h}\times Z_{h}\times W_{h}$ with
$\lvert\!\lvert\!\lvert(\boldsymbol{v},q,\tau,\xi)\rvert\!\rvert\!\rvert\lesssim\lvert\!\lvert\!\lvert(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})\rvert\!\rvert\!\rvert$,
@@
$\lvert\!\lvert\!\lvert(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})\rvert\!\rvert\!\rvert^{2}\lesssim B(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h};\boldsymbol{v},q,\tau,\xi).$
@@
$\lvert\!\lvert\!\lvert(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})\rvert\!\rvert\!\rvert^{2}$ $\lesssim B(\tilde{\boldsymbol{u}},\tilde{p},\tilde{\theta},\tilde{\lambda};\boldsymbol{v},q,\tau,\xi)-B(\boldsymbol{u}_{h},p_{h},\theta_{h},\lambda_{h};\boldsymbol{v},q,\tau,\xi)$
@@
(45) $\lesssim B(\tilde{\boldsymbol{u}}-\boldsymbol{u},\tilde{p}-p,\tilde{\theta}-\theta,\tilde{\lambda}-\lambda;\boldsymbol{v},q,\tau,\xi)+\mathcal{R}_{em},$
@@
where $\mathcal{R}_{em}:=\widetilde{a}_{h}(\boldsymbol{u},\boldsymbol{u},\boldsymbol{v})-\widetilde{a}_{h}(\boldsymbol{u}_{h},\boldsymbol{u}_{h},\boldsymbol{v})+\widetilde{c}_{h}(\boldsymbol{u},\theta,\tau)-\widetilde{c}_{h}(\boldsymbol{u}_{h},\theta_{h},\tau)+\langle\xi,g(\theta)-g(\theta_{h})\rangle_{\Gamma_{m}}$. Using the continuity bounds implies that
(46a) $|\widetilde{a}_{h}(\boldsymbol{u},\boldsymbol{u},\boldsymbol{v})-\widetilde{a}_{h}(\boldsymbol{u}_{h},\boldsymbol{u}_{h},\boldsymbol{v})|\leq M\big(\|(\boldsymbol{u}-\tilde{\boldsymbol{u}})\|_{1,h}+\|(\boldsymbol{u}_{h}-\tilde{\boldsymbol{u}})\|_{1,h}\big)\|\boldsymbol{v}\|_{1,h},$
(46b) $|\widetilde{c}(\boldsymbol{u},\theta,\tau)-\widetilde{c}(\boldsymbol{u},\theta,\tau)|\leq M\big(\|\theta-\tilde{\theta}\|_{1,\Omega}+\|\theta_{h}-\tilde{\theta}\|_{1,\Omega}+\|(\boldsymbol{u}-\tilde{\boldsymbol{u}})\|_{1,h}+\|(\boldsymbol{u}_{h}-\tilde{\boldsymbol{u}})\|_{1,h}\big)\|\tau\|_{1,\Omega},$
(46c) $|\langle\xi,g(\theta)-g(\theta_{h})\rangle_{\Gamma_{m}}|\leq L^{\prime}\|\xi\|_{-1/2,\Gamma_{m,h}}(\|\theta-\tilde{\theta}\|_{1,\Omega}+\|\theta_{h}-\tilde{\theta}\|_{1,\Omega}),$
where $M,L^{\prime}>0$ are sufficiently small. Combining (6.2) and (46) yields that
@@
$\lvert\!\lvert\!\lvert(\tilde{\boldsymbol{u}}-\boldsymbol{u}_{h},\tilde{p}-p_{h},\tilde{\theta}-\theta_{h},\tilde{\lambda}-\lambda_{h})\rvert\!\rvert\!\rvert^{2}\lesssim\lvert\!\lvert\!\lvert(\boldsymbol{u}-\tilde{\boldsymbol{u}},p-\tilde{p},\theta-\tilde{\theta},\lambda-\tilde{\lambda})\rvert\!\rvert\!\rvert^{2}+\sum_{K\in{\mathcal{T}}_{h}}h_{K}^{2}|\boldsymbol{u}-\tilde{\boldsymbol{u}}|_{2,K}^{2}.$
@@
Substituting the above bound in (6.2) leads to the stated estimate. Using the approximation results given in [22, 24] leads to the second stated result.
7 Lagrange multiplier stabilisation
We now briefly present a least-squares stabilised scheme in the case of boundary conditions associated with permeability or slip-type. In the conforming case, the discrete inf-sup condition to be satisfied is given by
However, in [38] it is shown that despite choosing stable inf-sup elements together with a typical choice for the Lagrange multiplier space as above, this condition may not be satisfied. To circumvent this difficulty, one can either enrich the velocity space with bubbles having compact support along $\Gamma_{\mathrm{m}}$ (see [38] for details), or add suitable residual stabilisation in the discrete problem (see, for example [39, 37]). We adopt the latter option. We define generic FE spaces $\boldsymbol{\mathrm{H}}_{h}\subset\boldsymbol{\mathrm{H}}$, $Q_{h}\subset Q$, $W_{h}\subset W$ and $Z_{h}\subset Z$ for the velocity, pressure, Lagrange multiplier, and concentration, respectively. Following [37], we first define the following mesh-dependent bilinear form $d_{h}:W_{h}\times W_{h}\to\mathbb{R}$:
(47) $d_{h}(\lambda_{h},\xi_{h}):=\sum_{e\in\mathcal{E}_{\mathrm{m}}}h_{e}\int_{e}\lambda_{h}\,\xi_{h}\,\mathrm{d}s,\qquad\forall\lambda_{h},\xi_{h}\in W_{h}.$
(48) $a(\boldsymbol{u}_{h},\boldsymbol{v}_{h})+\widetilde{a}(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{v}_{h})+b(\boldsymbol{v}_{h},(p_{h},\lambda_{h}))+s_{1}((\boldsymbol{u}_{h},p_{h},\lambda_{h}),\boldsymbol{v}_{h})$ $=0,$
Note that for the conforming method, the discrete quantities $\xi_{h},\,\boldsymbol{\sigma}_{h}\boldsymbol{n}\cdot\boldsymbol{n},\,\nabla\boldsymbol{v}_{h}\boldsymbol{n}\cdot\boldsymbol{n},\,q_{h}\mathbb{I}\boldsymbol{n}\cdot\boldsymbol{n}$ all belong to $W_{h}$. Note also that, for a Navier–Stokes model with slip boundary condition, [39] proved that choosing $\delta=0$ and $\alpha_{0}$ lower than a threshold yields a stable method. As in [20], for $\delta=1$ we have symmetry, however a smallness condition on $\alpha_{0}$ is needed for the sake of stability. For $\delta=-1$ we have the anti-symmetric variation of the method [16, 4], whose main advantage is the unconditional stability with respect to $\alpha_{0}$.
Following [37] we have that possible FE families which satisfy the inf-sup condition include Taylor–Hood ($\mathbb{P}_{2}-\mathbb{P}_{1}$), the MINI element ($\mathbb{P}_{1,b}-\mathbb{P}_{1}$), or Crouzeix–Raviart ($\mathbb{P}_{2,b}-\mathbb{P}_{1,\text{disc}}$) (see[13]). We remark that the $\boldsymbol{\mathrm{H}}^{1}(\Omega)$-nonconforming CR family can be used in this scheme without stabilisation. Indeed, it is enough to consider $\boldsymbol{\mathrm{H}}_{h}$ as in Section 4.1 together with $\alpha_{0}=0$. This choice reduces (48) to the discrete counterpart of (11), i.e., find $(\boldsymbol{u}_{h},p_{h},\lambda_{h},\theta_{h})\in\boldsymbol{\mathrm{H}}_{h}\times Q_{h}\times W_{h}\times Z_{h}$ such that
(49) $a_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})+\widetilde{a}_{h}(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{v}_{h})+b(\boldsymbol{v}_{h},(p_{h},\lambda_{h}))$ $=0,$
$b(\boldsymbol{u}_{h},(q_{h},\xi_{h}))$ $=\langle\xi_{h},g(\theta_{h})\rangle_{\Gamma_{\mathrm{m}}},$
for all $(\boldsymbol{v}_{h},q_{h},\xi_{h},\tau_{h})\in\boldsymbol{\mathrm{H}}_{h}\times Q_{h}\times W_{h}\times Z_{h}$ where $a_{h}(\cdot,\cdot)$ is the same as in Section 4.1.2, and
$\widetilde{a}_{h}(\boldsymbol{u}_{h};\boldsymbol{u}_{h},\boldsymbol{v}_{h}):=\rho_{0}\int_{\Omega}(\boldsymbol{w}\cdot\nabla_{h}\boldsymbol{u})\cdot\boldsymbol{v}.$
Consequently, one can follow the analysis in [24] to prove the well-posedness and convergence of this formulation.
@@ 8 Numerical experiments
We perform a series of computational tests using the finite element library FEniCS [1] together with the special module FeniCSii [26] for the treatment of bulk-surface coupling mechanisms. We perform an experimental error analysis through manufactured solutions. We monitor the errors of each individual unknown, the local convergence rate, and the number of necessary Newton–Raphson iterations to achieve convergence up to a prescribed tolerance of $10^{-7}$ on the residuals. By $e(\cdot)$ we denote the error associated with the quantity $\cdot$ in its natural norm, and denote by $h_{i}$ the mesh size corresponding to a refinement level $i$. The experimental convergence order is computed as
@@
To compute $\|\lambda-\lambda_{h}\|_{s,\Gamma_{\mathrm{m}}}$ (with $s\in\{-\frac{1}{2},\frac{1}{2}\}$ because we use Lagrange multipliers in these two spaces) we use the characterisation of $\mathrm{H}^{s}(\Gamma_{\mathrm{m}})$ in terms of the spectral decomposition of the Laplacian operator (see, e.g., [26]). For this, let $R:\mathrm{H}^{1}(\Gamma_{\mathrm{m}})\rightarrow\mathrm{H}^{1}(\Gamma_{\mathrm{m}})$ be the bounded linear operator defined by
@@
and so $\mathrm{H}^{s}(\Gamma_{\mathrm{m}})$ is the closure of the span of $\{r_{i}\}_{i=1}^{\infty}$ in this norm. During the experiments, different values for the stabilisation parameters are considered in order to capture the convergence of the method.
We also examine the behaviour of the schemes presented in Sections 4.1 and 7, focusing slightly more on the conservative scheme from Section 4.1. Furthermore, as we note in the experiments below, CR elements are more versatile as they can be used in both stabilised and non-stabilised schemes.
@@ 8.1 Divergence-conforming test
$k=2$
5071 0.141 1.24e-03 $\star$ 4.45e-03 $\star$ 6.50e-04 $\star$ 1.03e-05 $\star$ 7
19941 0.071 1.51e-04 3.04 5.75e-04 2.95 6.83e-05 3.25 1.28e-06 3.01 7
41703 0.049 4.91e-05 3.02 1.90e-04 2.98 2.11e-05 3.16 4.20e-07 3.00 7
75193 0.036 2.01e-05 3.01 7.84e-05 2.99 8.35e-06 3.12 1.73e-07 3.00 7
118483 0.029 1.01e-05 3.01 3.96e-05 2.99 4.12e-06 3.10 8.72e-08 3.00 7
177421 0.024 5.50e-06 3.01 2.16e-05 2.99 2.21e-06 3.08 4.75e-08 3.00 7
Table 8.1: Example 8.1. Error history and Newton iteration count for a FE family of $\mathbb{BDM}_{k+1}-\mathbb{P}_{k}-\mathbb{P}_{k}-\mathbb{P}_{k+1}$, with $k=0,1,2,$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively, on the unit square domain $\Omega=(0,1)^{2}$. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$-norm, and $\alpha_{0}=10(k+2)$.
@@
$\boldsymbol{u}(x,y,z)=\left(\begin{aligned} &\sin(\pi z)\,\cos(\pi y)\\ &-\cos(\pi x)\,\sin(\pi z)\\ &\sin(\pi x)\,\cos(\pi y)\end{aligned}\right),\quad p(x,y,z)=\sin(x^{2}+y^{2}+z^{2}),\quad\hskip-2.84544pt\theta(x,y)=e^{-xyz}.$
Here we observe that $\boldsymbol{u}$ is solenoidal, and again we consider $D_{0}=\rho_{0}=\mu_{0}=1$. We choose $k\in\{0,1\}$ in order to study the convergence rates on different polynomial orders. Table 8.2 present the error history, mesh sizes and number of iterations for the stabilisation parameter $\alpha=10(k+2)$. Here, an optimal $\mathcal{O}(h^{k+1})$ convergence order is observed for k=0, 1k=0,\,1.
@@
222 1.000 2.32e+$0$ $\star$ 1.73e+$0$ $\star$ 2.81e+$0$ $\star$ 1.80e-01 $\star$ 4
4478 0.554 1.11e+$0$ 1.26 1.04e+$0$ 0.87 9.22e-01 1.89 8.82e-02 1.21 4
@@
Table 8.2: Example 8.1. Error history and Newton iteration count for a FE family of $\mathbb{BDM}_{k+1}-\mathbb{P}_{k}-\mathbb{P}_{k}-\mathbb{P}_{k+1}$, with $k=0,1,$ for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively, on the unit cube domain $\Omega=(0,1)^{3}$. For this case the Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$-norm, and $\alpha_{0}=10(k+2)$.
Finally, we show the results obtained with the $\mathbb{CR}_{1}-\mathbb{P}_{0}$ pair for velocity-pressure instead of the $\mathbb{BDM}_{1}-\mathbb{P}_{0}$ pair. We consider the same manufactured solution as before and we test the scheme in two and three dimensions. We recall that piecewise constants are used to approximate the Lagrange multiplier. Table 8.3 describes the behaviour of the scheme with a stabilisation parameter $\alpha_{0}=20$, indicating a similar accuracy as in the $\boldsymbol{\mathrm{H}}^{1}(\Omega)$-conforming scheme presented in Table 8.1.
Stabilisation using $\mathbb{CR}_{1}-\mathbb{P}_{0},\,\alpha_{0}=20$, $\Omega=(0,1)^{2}$
51 0.707 1.63e+$0$ $\star$ 1.92e+$0$ $\star$ 1.97e+$0$ $\star$ 1.87e-01 $\star$ 6
103 0.471 1.15e+$0$ 0.87 1.89e+$0$ 0.05 1.28e+$0$ 1.08 1.19e-01 1.11 6
261 0.283 7.08e-01 0.95 1.38e+$0$ 0.62 5.60e-01 1.61 7.15e-02 1.00 7
793 0.157 3.91e-01 1.01 8.52e-01 0.82 1.90e-01 1.84 4.00e-02 0.99 7
2721 0.083 2.05e-01 1.02 4.77e-01 0.91 5.56e-02 1.93 2.12e-02 0.99 7
10033 0.043 1.05e-01 1.01 2.52e-01 0.96 1.50e-02 1.97 1.10e-02 1.00 7
Stabilisation using $\mathbb{CR}_{1}-\mathbb{P}_{0},\,\alpha_{0}=20$, $\Omega=(0,1)^{3}$
222 1.000 2.29e+$0$ $\star$ 2.37e+$0$ $\star$ 2.65e+$0$ $\star$ 1.85e-01 $\star$ 4
4478 0.554 1.02e+$0$ 1.37 8.10e-01 1.82 5.10e-01 2.79 8.84e-02 1.26 4
22466 0.273 5.48e-01 0.88 4.45e-01 0.85 2.14e-01 1.23 4.70e-02 0.89 4
139391 0.137 2.80e-01 0.97 2.26e-01 0.98 6.90e-02 1.63 2.44e-02 0.95 4
256159 0.112 2.25e-01 1.11 1.80e-01 1.15 4.86e-02 1.78 1.96e-02 1.10 4
Table 8.3: Example 8.1. Error history for $\mathbb{CR}_{1}-\mathbb{P}_{0}-\mathbb{P}_{0}-\mathbb{P}_{1}$ approximations for $\boldsymbol{u}_{h},p_{h},\lambda_{h}$ and $\theta_{h}$, respectively. The Lagrange multiplier errors are measured in the $\mathrm{H}^{-1/2}(\Gamma_{\mathrm{m}})$-norm and we considered two and three dimensional domains.
8.2 Lagrange stabilisation test